OpenStudy (anonymous):

So I am transforming parent functions to other certain equations that apply to the description, and I'm not sure how to write them out. For example, y= the square root of X, Translate 5 units down; Vertical stretch by a factor of 3/2.

OpenStudy (anonymous):

What do you know to do so far? Any guesses? Parent function is.........

OpenStudy (anonymous):

Well for the equation, it's a no go, but the parent funtion is the main function you go off of, like in my example \[ y= \sqrt{x} is the parent function.\]

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

I'm sure you know what I means. Haha

OpenStudy (anonymous):

y = sqrt(x) Vertical shifts occur when we add/subtract something after the function Vertical stretches/shrinks occur when we mult something time the function. Horizontal shifts occur when we add/subtract from the variable What might you guess now?

OpenStudy (anonymous):

I might guess the point on the graph would be pulled down to -5, but still a bit foggy with the vertical stretch.

OpenStudy (anonymous):

You have the idea about what happens with the shift. All the point of the parent graph move down 5 to create the new one. y =sqrt(x) -5 The stretch/shrink literally makes every value of the original function larger (stretch) or smaller (shrink). This happens when we mult something times the function. so....

OpenStudy (anonymous):

The factor 3/2 \[\frac{ 3 }{ 2 } \times -5\]

OpenStudy (anonymous):

Or possibly the other way around.

OpenStudy (anonymous):

So close. y = (3/2) sqrt(x) -5 ^ ^ ^ vert stretch/shrink function vertical shift

OpenStudy (anonymous):

The trick is being able to do this with a different situation.